# The EU AI Act, Stakeholder Needs, and Explainable AI: Aligning Regulatory Compliance in a Clinical Decision Support System

### v2

@@ 3 Methodology
To answer our research question, we apply a two-pronged methodology within a cross-functional team. This section first describes the team and the distribution of responsibilities and incentives. We then describe the CDSS. That is followed by the two prongs of the approach, first the legal analysis of the AI Act and then the XAI analysis, both in relation to the CDSS. We then show how the two prongs merge.
3.1 The Cross-functional Team
The team behind this contribution consists of five individuals. Two individuals represent a company developing the CDSS. One of them has a PhD in computer science and has extensive experience in requirements engineering and empirical software engineering. The other is a PhD student who focuses on the human-centered implementation of XAI methods for AI-based CDSSs. Human-centered XAI methods put the human and their desiderata at the center of the implementation of XAI methods. The daily work of these individuals takes place in the context of an interdisciplinary exchange with the clinicians in a publicly funded research project. They organized the workshops (see Section˜4) with usability engineers from the company, but they are not part of the stakeholders that have been analyzed.
Two individuals represent a research institute; one with a background as a judge at an appeals court, as well as working for multiple governmental investigations regarding the introduction of new laws, and the other with a PhD in computer science with a focus on model-driven engineering and natural language processing. Together, they have assisted the Government’s Office during the negotiation and implementation of the AI Act and have shared their insights with the EU Commission. They also assist a notified body in defining the implications of the AI Act to existing and possible new business opportunities, e.g., when a notified body is needed, how to show fulfillment within a conformity assessment procedure, and to which concrete products the AI Act is applicable.
The fifth member of the author team is a university professor for information systems and AI. The primary responsibility in relation to this contribution is to serve as an external auditor, assessing the validity and scientific soundness of the contribution and delivering both a technical and socio-technical viewpoint.
@@ 3.2 The Clinical Decision Support System
This is where the CDSS comes into play. It supports clinicians by predicting key indicators that guide the transfer decision—such as remaining LOS, mortality probability, and readmission probability—using vital signs like heart rate and oxygen saturation from the past 24 hours. In particular, when the AI’s predictions conflict with a clinician’s intuition, the interpretability of the system is crucial for reconsidering and validating the final decision. To generate these predictions, a DNN is trained on time-series patient data spanning the years 2017 to 2024. Notably, the system does not suggest which patients to move, but rather shows the predictions to help clinicians make their own decisions.
@@ 3.3 Legal Analysis
The legal analysis consists of mapping the different articles of the AI Act to the system under investigation—which technical and organizational requirements does the system imply? Then, who will be responsible for the fulfillment of the requirements? The details of the technical requirements are not as important for the analysis as their implications for the different roles. For the legal analysis, e.g., it is more important to know who is responsible for which aspects of data management (input data, validation data, etc.), than how to assess the relevance or completeness of data. As analyses emerge, we visualize the concepts and their relationships since text has the drawback of enforcing a linear representation (see e.g. Figure˜2(a)). The primary outcome of the legal analysis is given in Section˜4.
At the same time, we make decisions on what is in and out of scope. The latter is necessary since the AI Act consists of 113 articles that define roles, responsibilities, technical requirements, fines, and so on. That is complemented by 13 annexes that supply further detail to the articles. To motivate and help with the interpretation of the articles and the annexes, there are 180 recitals. Knowing what to include in an analysis is therefore an analysis in itself and something that is made easier as more is known about the system under consideration. In the case of this contribution, we settle on focusing primarily on the responsibilities of the roles of the provider and deployer, leaving out the content related to what happens if a system does not comply. We also keep the level of detail abstract in terms of technical requirements as going into detail would yield enough content for multiple academic contributions per article of the AI Act.
In parallel, we conduct analyses of the AI Act in relation to the new machinery regulation as an assignment for the notified body (c) and a first analysis of the cybersecurity requirements of the AI Act in relation to the regulation on medical devices and the Cyber Resilience Act (anonymized). These parallel investigations also suffice to better understand the AI Act in relation to the CDSS.
@@ 3.4 Requirements Engineering Activities
To understand the requirements of the clinicians who are going to use the CDSS, we conduct a workshop in which we collect the most important functional and quality requirements. The workshop is modeled after the Design Thinking method (ferreira2019design), but much abbreviated. Participants were clinicians and ML researchers, with a requirements engineer moderating the workshop.
@@ 3.5 Stakeholder Mapping and XAI Analysis
The next step of analysis is to identify the relevant stakeholder groups. Thereby, we have discussions involving the AI Act and XAI experts, following the stakeholder mapping approach outlined by langer_what_2021. Stakeholders identified during the Design Thinking Workshop are mapped, and we summarize their specific desiderata and expand upon them in regular project meetings and exchanges. For the choice of XAI methods, the categorization by speith_review_2022 is used. The Ph.D. student analyzes the problem and selects common methods that align with the problem types and relevant requirements of the CDSS, as identified in the requirements engineering workshops.
@@ 3.6 The Merging of the two Approaches
The legal and the XAI analysis are made in parallel so that they benefit from insights from the other approach. This takes place over three different phases. Phase one was a one-day sprint (March 2024) where the AI Act is introduced to the company and a subsequent workshop starts to map out the high-level applicability towards the CDSS. Phase two occurs six months later (September 2024) to share new insights in a second workshop. The third phase runs from December 2024 to January 2025 and is conducted as six weekly meetings with the ambition to document the combined analysis and the reasoning. Between the meetings of the third phase, each individual contributes to the shared document, and the meetings discuss the combined effort, and changes are made as new insights are gained. The outcome in this contribution represents the final and polished analyses.
@@ 3.7 Validity Strategies
Following LincolnGuba1985, we structure the validity strategies along Credibility (establishing an analysis true to the data and context); Transferability (ensuring that the findings were applicable beyond the context); Dependability (enabling repetition of our contribution); and Confirmability (limiting the impact of our own incentives and bias).
@@ 4 Applicability of the AI Act to the CDSS
We analyzed the CDSS in the context of the EU AI Act and establish whether the act is applicable to the case of the CDSS system. To do this, we need to answer three questions:
@@
  • Does the system need to be CE-marked? (Art. 6.)
Scope: The AI Act has a number of clauses defining the scope of the regulation. For instance, it “does not apply to any research, testing or development activity regarding AI systems or AI models prior to their being placed on the market or put into service” or systems used in a purely personal, non-professional activity. After going through the scope as defined in Art. 2, our assessment is that the AI system under investigation is within scope of the AI Act as it is to be put into use in a clinical context.
AI system: An AI system in the AI Act is defined as “a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments”.
A key concept here is infer, which can be done by machine-learning approaches but also through knowledge- or logic-based technologies. However, “systems that are based on the rules defined solely by natural persons to automatically execute operations” should not be seen as AI systems (recital 12). Our assessment is that the AI system under investigation is within the scope of the definition of an AI system as it, from received input, generates a prediction that, if the clinician approves, will influence where and by whom the patient will be treated.
CE-mark: There are two scenarios where the AI system needs to be CE-marked: the first covers clinical systems that fall under EU regulation 2017/745 on medical devices and subsequently need to undergo a conformity assessment procedure involving a notified body (Art. 6.1 of the AI Act); the second is for AI systems that are listed in Annex III of the AI Act. In the latter case, point 5.a refers to “AI systems intended […] to evaluate the eligibility of natural persons for […] healthcare services, as well as to grant, reduce, revoke, or reclaim such […] services”.
@@
Taken together, this implies that the AI system will need to be CE-marked through internal compliance, i.e., without involving a notified body (Annex IV of the AI Act).
Conclusion: The AI system needs to be CE-marked through an internal conformity assessment procedure in accordance with the AI Act. This is the obligation of the provider. In practice, this means that the provider needs to fulfill several obligations, as described in Section˜2.1. This includes fulfilling technical criteria (Art. 9-15) and having a quality management system (Art. 17).
That is mirrored by the obligations of the deployer (Art. 26), also described in Section˜2.1. Since the system will be deployed by a legal person governed by public law (or by private entities providing public services), there is another obligation: a Fundamental Rights Impact Assessment shall be conducted by the deployer (Art. 27). How to do this exactly will be further detailed in a harmonized standard.
@@ 5 Choosing the right XAI Method
Choosing the right XAI method is important to ensure that the stakeholder needs are satisfied (schoonderwoerd_human-centered_2021). Therefore, we analyze relevant stakeholders and elaborate their desiderata (see Section˜5.1). This information helps us make well-founded choices about the XAI method in the implementation phase of the CDSS (see Section˜5.2).
@@ 5.1 Our Stakeholders and their Desiderata
The first group of stakeholders are researchers and engineers specializing in ML, who are categorized as developers and identified as providers under the provisions of the AI Act. Their primary requirement for the AI system is the achievement of high performance and reliability. In order to fulfill this requirement, they concentrate on quality measures such as accuracy and robustness, which are of the highest importance for the practical implementation and success of the model in real-world applications. This will increase end-users’ acceptance of our AI system, which is a key factor for hospital management.
@@
Clinicians are the primary users of the AI system in CDSS, as they are the only individuals who directly interact with the AI outputs and ultimately make the final decisions. Most of these clinicians are novice users; as defined in Section˜2.2, the AI system must be accessible and user-friendly. This translates into the desideratum interpretability, which involves providing comprehensible, plausible, and correct explanations for the AI system. This interpretability allows clinicians to question AI predictions rather than simply accepting them, thereby fostering trust and appropriate reliance (schemmer_appropriate_2023). Clinicians should trust the AI system when it is correct and appropriately rely on their judgment when the AI system is incorrect. This is most important when the AI system makes a false prediction that negatively impacts the patient, e.g., the AI system falsely predicts a short remaining LOS for the patients, although the patient is very sick and will have to stay longer. In this situation, the clinicians must rely on their medical expertise.
Patients, although not directly interacting with the AI system, are significantly impacted by the decisions made by clinicians using the AI system and thus an “affected party” according to the definition in Section˜2.2. Patients have a strong need for correct and fair decision-making. Fairness in AI-driven decisions is a critical requirement, as it directly affects patient outcomes and trust in the healthcare system.
The regulator corresponds in our case to the competent authority, which can be mapped to different AI Act stakeholder roles, e.g., AI Office or AI Board.
@@ 5.2 Our XAI Method Choices
After we clarified the relevant stakeholders and their desiderata, we now describe the concrete XAI method choices we made for our CDSS. We made choices in the categories applicability, scope, and stage, as proposed by speith_review_2022. Categories like input data and problem type are pre-determined by the problem the application addresses.
Input Data, Problem, and Model Type: The AI system is based on a complex DNN that processes multivariate time series data, including patient vital signs, to predict the remaining LOS of an individual patient.
Stage: While DNNs provide great performance by learning complex relationships in data, they are not intrinsically interpretable and lack transparency (barnes_evaluating_2024). Therefore, we employ post-hoc XAI methods to provide insights into this black-box model.
Applicability: To ensure a reliant and permanent AI model, we dynamically retrain and update the AI model to avoid data drifts. Some changes in the AI model might affect the AI architecture. Therefore, the model-specific XAI method could not be applied anymore. For this reason, we selected model-agnostic XAI methods. This choice allows us to apply the XAI method to any underlying AI model that is based on a neural network structure.
Scope: The primary goal of our XAI method is to satisfy our end user’s desiderata and thus target interpretability, usability, and trust or appropriate reliance. For this reason, our XAI method focuses on a local scope, enabling clinicians to understand specific predictions. This local scope is done by interpreting relationships between input data, such as patient vital signs, and predicted outcomes, like the estimated remaining LOS. This model’s interpretability is crucial to ensure an appropriate reliance on clinical decision-making.
Output Format and Result: To be sure that the output of the XAI method is understood accordingly, we are using a combination of different output formats for the XAI method result, such as visual and numerical information. This degree of freedom allows us to provide an appropriate output format for users with different desiderata and levels of expertise.
Figure 1: Example Heatmap Explanation with Shap. Shows the five most important features in descending order of relevance over a period of 14 hours. The color gradient indicates whether the feature value at a given timestamp has a positive (red) or negative (positive) impact on the predicted LOS.
Finally, the different choices in the categories allow us to choose between XAI methods such as Shap (lundberg_unified_2017), LIME (ribeiro_why_2016), Integrated Gradient (sundararajan_axiomatic_2017), or Saliency Maps (simonyan_deep_2014). Functionally-grounded evaluation methods (doshi-velez_towards_2017) that measure explanation properties such as complexity, faithfulness, or sensitivity help us to select the best XAI method (chenWhatMakesGood2024).
@@ 6 Contrasting XAI with the AI Act
From analyzing the same CDSS from the perspectives of XAI and the AI Act, we have created opportunities to reflect on what one perspective reveals about the other, thus a way to answer our research question.
The AI Act refers to stakeholder obligations, while XAI describes stakeholder desiderata. Mapping these two perspectives is not a straightforward one-to-one mapping. Rather than exhaustively detailing the inter-relationships, we explored a thin slice of their intersection to show how XAI can fulfill the AI Act’s legal requirements. This means we are explicitly expecting further research to complement and refute our findings (see KI1-KI3), both in scope and in conclusions, if XAI is chosen as a relevant way to comply with the requirements of the AI Act. A schematic view of the analysis is found in diagrams of Figure˜2.

### v3

@@ 3 Methodology
To answer our research question, we apply a two-pronged methodology within an interdisciplinary team. This section first describes the team and the distribution of responsibilities and incentives. We then describe the CDSS. That is followed by the two prongs of the approach: the legal analysis of the AI Act and the XAI analysis, both in relation to the CDSS. We then show how the two prongs merge.
3.1 The Interdisciplinary Team
The team behind this contribution consists of five individuals representing diverse expertise. Two individuals from the company developing the CDSS bring backgrounds in requirements engineering, empirical software engineering, and human-centered XAI implementation. They organized the stakeholder workshops (see Section˜4) and maintain ongoing interdisciplinary exchange with clinicians within a publicly funded research project. Two individuals from a research institute contributed legal and technical AI Act expertise: one with experience as an appeals court judge and involvement in governmental legislative investigations, the other with a PhD in computer science. Together, they assisted the Government’s Office during AI Act negotiations and implementation, and currently advise notified bodies on conformity assessment procedures and AI Act applicability. The fifth member is a university professor for information systems and AI, serving as an external auditor to assess the validity and scientific soundness of the contribution from both technical and socio-technical perspectives.
@@ 3.2 The Clinical Decision Support System
This is where the CDSS comes into play. It supports clinicians by predicting key indicators that guide the transfer decision — such as remaining LOS, mortality probability, and readmission probability — using vital signs like heart rate and oxygen saturation as well as indicators such as the need for ventilation or the presence of an infection. When the AI’s predictions conflict with a clinician’s intuition, the interpretability of the system is crucial for reconsidering and validating the final decision. To generate these predictions, a deep neural network (DNN) is trained on time-series patient data spanning the years 2017 to 2024. Notably, the system does not suggest which patients to move, but rather shows the predictions to help clinicians make their own decisions.
@@ 3.3 Legal Analysis
The legal analysis consists of mapping the different articles of the AI Act to the system under investigation. Which technical and organizational requirements does the system imply? Who will be responsible for the fulfillment of the requirements? The details of the technical requirements are not as important for the analysis as their implications for the different roles. For the legal analysis, e.g., it is more important to know who is responsible for which aspects of data management (input data, validation data, etc.), than how to assess the relevance or completeness of data. As analyses emerge, we visualize the concepts and their relationships since text has the drawback of enforcing a linear representation (see e.g. Figure˜2(a)). The primary outcome of the legal analysis is given in Section˜4.
At the same time, we make decisions on what is in and out of scope. This is necessary since the AI Act consists of 113 articles that define roles, responsibilities, technical requirements, and so on. That is complemented by 13 annexes that supply further detail to the articles. To motivate and help with the interpretation of the articles and the annexes, there are 180 recitals. Knowing what to include in an analysis is therefore an analysis in itself and something that is made easier as more is known about the system under consideration. In the case of this contribution, we settle on focusing primarily on the responsibilities of the roles of the provider and deployer, leaving out the content related to what happens if a system does not comply. We also keep the level of detail abstract in terms of technical requirements as going into detail would yield enough content for multiple academic contributions per article of the AI Act.
@@ 3.4 Requirements Engineering Activities
To understand the requirements of the clinicians who are going to use the CDSS, we conduct a workshop in which we collect the most important functional and quality requirements. The workshop is modeled after the Design Thinking method (Ferreira Martins et al., 2019), but much abbreviated. Participants were clinicians and ML researchers, with a requirements engineer moderating the workshop.
@@ 3.5 Stakeholder Mapping and XAI Analysis
The next step of analysis is to identify the relevant stakeholder groups. We follow the stakeholder mapping approach outlined by Langer et al. (2021) and involve the AI Act and XAI experts. Stakeholders identified during the Design Thinking Workshop are mapped, we summarize their specific desiderata, and expand on them in regular project meetings and exchanges. For the choice of XAI methods, the categorization by Speith (2022) is used. The XAI expert analyzes the problem and selects common methods that align with the problem types and relevant requirements of the CDSS, as identified in the requirements engineering workshops.
@@ 3.6 The Merging of the two Approaches
The legal and the XAI analysis are conducted in parallel so that they benefit from insights from the other approach. This takes place over three different phases. Phase one is a one-day sprint (March 2024) where the AI Act is introduced to the company and a subsequent workshop starts to map out the high-level applicability towards the CDSS. Phase two occurs six months later (September 2024) to share new insights in a second workshop. The third phase runs from December 2024 to January 2025 and is conducted as six weekly meetings with the ambition to document the combined analysis and the reasoning. Between the meetings of the third phase, each individual contributes to the shared document, the meetings discuss the combined effort, and changes are made as new insights are gained. The outcome in this contribution represents the final analyses.
@@ 3.7 Validity Strategies
Following Lincoln and Guba (Lincoln and Guba, 1985), we structure the validity strategies along Credibility (establishing an analysis true to the data and context); Transferability (ensuring that the findings were applicable beyond the context); Dependability (enabling repetition of our contribution); and Confirmability (limiting the impact of our own incentives and bias).
@@ 4 Applicability of the AI Act to the CDSS
As the first step of our overall analysis and as described in Section˜3.3, we analyze the CDSS in the context of the EU AI Act and establish whether the act is applicable to the case of the CDSS system. To do this, we need to answer three questions:
@@
  • Does the system need to be CE-marked? (Art. 6)
Scope.
The AI Act has a number of clauses defining the scope of the regulation. For instance, it “does not apply to any research, testing or development activity regarding AI systems or AI models prior to their being placed on the market or put into service” or systems used in a purely personal, non-professional activity. After going through the scope as defined in Art. 2, our assessment is that the AI system under investigation is within scope of the AI Act as it is to be put into use in a clinical context.
AI system.
An AI system in the AI Act is defined as “a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments”. A key concept is infer, which can be done by machine-learning approaches but through knowledge- or logic-based technologies. However, “systems that are based on the rules defined solely by natural persons to automatically execute operations” should not be seen as AI systems (recital 12). Our assessment is that the AI system investigated is within the scope of the definition of an AI system as it, from received input, generates a prediction that, if the clinician approves, will influence where and by whom the patient will be treated.
CE-mark.
There are two scenarios where the AI system needs to be CE-marked: the first covers clinical systems that fall under EU regulation 2017/745 on medical devices and subsequently need to undergo a conformity assessment procedure involving a notified body (Art. 6.1 of the AI Act); the second is for AI systems that are listed in Annex III of the AI Act. In the latter case, point 5.a refers to “AI systems intended […] to evaluate the eligibility of natural persons for […] healthcare services, as well as to grant, reduce, revoke, or reclaim such […] services”.
@@
To summarize, the AI system needs to be CE-marked through an internal conformity assessment procedure in accordance with the AI Act, i.e., without involving a notified body (Annex IV of the AI Act). This is the obligation of the provider. In practice, this means that the provider needs to fulfill several obligations, as described in Section˜2.1. This includes fulfilling technical criteria (Art. 9-15) and having a quality management system (Art. 17).
That is mirrored by the obligations of the deployer (Art. 26), described in Section˜2.1. Since the system will be deployed by a legal person governed by public law (or by private entities providing public services), there is another obligation: a fundamental rights impact assessment (FRIA) shall be conducted by the deployer (Art. 27). How to do this will be detailed in a harmonized standard.
We use our findings from this legal analysis in Section˜6 as the foundation for contrasting XAI and the AI Act.
@@ 5 Choosing the right XAI Method
Choosing the right XAI method is important to ensure that stakeholder needs are satisfied (Schoonderwoerd et al., 2021). Therefore, we analyze relevant stakeholders and elaborate their desiderata (see Section˜5.1). This information helps us make well-founded choices about the XAI method in the implementation of the CDSS (see Section˜5.2).
@@ 5.1 Our Stakeholders and their Desiderata
The first group of stakeholders are researchers and engineers specializing in machine learning (ML), who are categorized as developers and identified as providers under the provisions of the AI Act. Their primary requirement for the AI system is the achievement of high performance and reliability. In order to fulfill this requirement, they concentrate on quality measures such as accuracy and robustness, which are of the highest importance for the practical implementation and success of the model in real-world applications. This will increase end-users’ acceptance of our AI system, which is a key factor for hospital management.
@@
Clinicians are the primary users of the AI system in CDSS, as they are the only individuals who directly interact with the AI outputs and ultimately make the final decisions. Most of these clinicians are novice users; as defined in Section˜2.2, the AI system must be accessible and user-friendly. This translates into the desideratum interpretability, which involves providing comprehensible, plausible, and correct explanations for the AI system. This interpretability allows clinicians to question AI predictions rather than simply accepting them, thereby fostering trust and appropriate reliance (Schemmer et al., 2023). Clinicians should trust the AI system when it is correct and appropriately rely on their judgment when the AI system is incorrect. This is most important when the AI system makes a false prediction that negatively impacts the patient, e.g., the AI system falsely predicts a short remaining LOS for the patients, although the patient is very sick and will have to stay longer. In this situation, the clinicians must rely on their medical expertise.
Patients, although not directly interacting with the AI system, are significantly impacted by the decisions made by clinicians using the AI system and thus an “affected party” according to the definition in Section˜2.2. Patients have a strong need for correct and fair decision-making. Fairness in AI-driven decisions is a critical requirement, as it directly affects patient outcomes and trust in the healthcare system. The regulator corresponds to the competent authority, which can be mapped to different AI Act stakeholder roles, e.g., AI Office or AI Board.
@@ 5.2 Our XAI Method Choices
After we clarified the relevant stakeholders and their desiderata, we describe the concrete XAI method choices we made for our CDSS. Categories like input data and problem type are pre-determined by the problem the application addresses.
The AI system is based on a complex DNN that processes multivariate time series data, including patient vital signs, to predict the remaining LOS of an individual patient. While DNNs provide great performance by learning complex relationships in data, they are not intrinsically interpretable and lack transparency (Barnes and Hutson, 2024). Therefore, we employ post-hoc XAI methods to provide insights into this black-box model. To ensure a reliable and permanent AI model, we dynamically retrain and update the AI model to avoid data drifts. Some changes in the AI model might affect the AI architecture, making model-specific XAI methods potentially incompatible. For this reason, we selected model-agnostic XAI methods that can be applied to any underlying AI model based on a neural network structure.
The primary goal of our XAI method is to satisfy our end users’ desiderata and thus target interpretability, usability, and trust or appropriate reliance. Our XAI method focuses on a local scope, enabling clinicians to understand specific predictions by interpreting relationships between input data, such as patient vital signs, and predicted outcomes, like the estimated remaining LOS. This interpretability is crucial to ensure appropriate reliance in clinical decision-making. To ensure that the output of the XAI method is understood accordingly, we use a combination of different output formats, such as visual and numerical information, providing appropriate representations for users with different desiderata and levels of expertise.
Figure 1: Example Heatmap Explanation with SHAP. Shows the five most important features in descending order of relevance over a period of 14 hours. The color gradient indicates whether the feature value at a given timestamp has a positive (red) or negative (green) impact on the predicted LOS.
Finally, the different choices in the categories allow us to choose between XAI methods such as SHAP (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016), Integrated Gradient (Sundararajan et al., 2017), or Saliency Maps (Simonyan et al., 2014). Functionally-grounded evaluation methods (Doshi-Velez and Kim, 2017; Nauta et al., 2023) that measure explanation properties such as complexity, faithfulness, or sensitivity help us to select the best XAI method (Chen et al., 2024).
@@ 6 Contrasting XAI with the AI Act
By analyzing the same CDSS from the perspectives of XAI and the AI Act, we have created opportunities to reflect on what one perspective reveals about the other. This allows us to answer our research question.
The AI Act refers to stakeholder obligations and responsibilities, while human-centered XAI research focuses on stakeholder desiderata. Mapping these two perspectives is not a straightforward one-to-one mapping. Rather than exhaustively detailing the inter-relationships, we explored a thin slice of their intersection to show how XAI can fulfill the AI Act’s legal requirements. Our main finding is that the user-centered implementation of XAI enables conformant representation. A schematic view of the analysis is found in diagrams of Figure˜2. This means we are explicitly expecting further research to complement and refute our findings (see KI1-KI3), both in scope and in conclusions, if XAI is chosen as a relevant way to comply with the requirements of the AI Act.